Table 2 shows the frequency distribution of male cases and controls by average number of cigarettes smoked per day.

Table 2. Most recent amount of cigarettes smoked daily before onset of the present illness, lung cancer cases and matched controls with other diseases, Great Britain, 1948-1952.

Daily number of cigarettes # Cases # Controls Odds Ratio

0

8

68

referent

1-14

622

785

15-24

498

482

25+

402

195

All smokers

1,522

1,462

Total

1,530

1,530


***Question 11: Compute the odds ratio by category of daily cigarette consumption, comparing each smoking category to nonsmokers??

#11. HINT You'll end up with 4 Odds Ratios on this one. One for each dosage of cigarettes (1-14; 15-24; 25+; and ALL). These are NOT age categories; they are dosage (# of cigarettes smoked/day). Set up your 2x2 table for each category. Your 'no exposure' row (0 cigarettes) will be the same for all the 2x2 tables.

*****Question 12: From Question 11; what does these results tell you? HINT: Look at the 4 different ORs and note any differences by dosage. Is there a dose-response relationship between exposure (# of cigarettes smoked) and disease (lung cancer)?
Answer #1

Question 11:- odds ratio for each category is as below

Coe contos G8 2. 622 4 18 02 S-24 ts ,a 1,2 ,s30 s30 1,462 Totalodds ratio = ad/bc

Odds vaho hy catepoy of condumphon, compang each smo Cate n9 gooy to non-smokens f we tate Seatal lumbesuotse rouor g ta) 6&(odds ratio= 8×785/68×622

= 0.15

o ( non-Smo tuu) 3 et fad ) 498 (C)odds ratio = 8×482/ 68×498

=0.11

on-smo astodds ratio = 8×195/68×402

=0.06.

S. ,S2 1,96odds ratio= 8×1462/68×1522

=0.11

Question 12:-

Odds ratio for each category is as follows:-

Number of cigarettes Odds ratio
1-14 0.15
15 - 24 0.11
25+ 0.06
All 0.11

Results:- number of cigarettes smoked directly relates to the occurrence of lung cancer